A U.S. firm has a debt obligation of ¥ 194 million payable in one year. The current spot rate is ¥ 116 per U.S. dollar and the one-year forward rate is ¥ 110 per U.S. dollar. Additionally, a one-year Call option on the Yen with a strike price of $0.0085 per yen can be purchased for a premium of 0.011 cent per yen. The risk-free money-market rate in Japan is 1.6% and the risk-free money-market rate in the U.S. is 3.9%. Calculate the future U.S. dollar cost of meeting this obligation using a call option hedge. Assume that at expiration, the spot rate will be above the option strike price such that the option contract will be exercised. Don't forget to adjust the Call premium payment made at the beginning of the year by the time value of money since you are calculating the future dollar cost, that is, at the end of the year.
Answer #1

Premium Payable TODAY = Yen Required*Premium per Yen = 194000000*0.011cents = 2134000Cents = $21340.00

Future Value of Premium Paid = Premium*(1+Interest Rate) = 21340*[1+0.039] = $22172.26

US Dollar Cost of meeting obligation(if Options Hedge is used) = US Dollars Required to Buy the Yen under Options Contract + Future Value of Premium Paid

= [0.0085*194000000] + 22172.26

= $1671172.26